OpenStudy (anonymous):

What mass of lead (density 11.4 g/cm3) would have an identical volume to 17.7 g of mercury (density 13.6 g/cm3)?

OpenStudy (anonymous):

the given mercury slightly exceeds 1 cm^3 in volume. 13.6g would take space of 1 cm1, we have 17.7g. \[\frac{ 17.7g }{x cm^{3}}=\frac{13.6g}{1 cm^{3}}\] \[\frac{ x cm^{3} }{17.7g}=\frac{1 cm^{3}}{13.6g}\] \[x cm^{3}=\frac{17.7 g cm^{3}}{13.6g}\] 17.7g of mercury has 17.7/13.6 = 1,3 cm^3 in volume.

OpenStudy (anonymous):

since every cm^3 of lead has 11.4 g, and we want the identical volume of 1,3 cm^3: 1,3 cm^3 * 14.4g/cm^3 =14.4 g x 1.3 = 18.72 g